Find the LCM of the following Polynomials:
x^3 - x and x^2 - x
Answer by stanbon(75887) 
 Find the LCM of the following Polynomials:
x^3 - x and x^2 - x
-------
Factor each expression:::
x^3-x = x(x^2-1) = x(x-1)(x+1)
------
x^2-x = x(x-1)
-------
LCM = x(x-1)
=================
Cheers,
Stan H.
=================
|
|
|